OpenStudy (anonymous):

1. What is the degree of hte monomial 15x^3y^9? A) 12 B) 3 C) 9 D) 15 2. What is the difference? 10c^3 - 8c^3 A) 18c^3 B) 2c^3 C) 3c^6 D) -80c^6 3. What is the sum? 16x^3 + 9x^3 25x^6 25x^3 7x^3 24x^6

Vocaloid (vocaloid):

lets start with #1 add the exponents together

OpenStudy (anonymous):

@Vocaloid 12

Vocaloid (vocaloid):

good, so thats the degree

Vocaloid (vocaloid):

ready for #2?

OpenStudy (anonymous):

Yes

Vocaloid (vocaloid):

ok, this is just combining like terms if 10 - 8 = 2, then what is 10c^3 - 8c^3 = ?

OpenStudy (anonymous):

@Vocaloid 2c^3

Vocaloid (vocaloid):

good!

Vocaloid (vocaloid):

ready for the next one?

OpenStudy (anonymous):

Yes

Vocaloid (vocaloid):

16 + 9 = 25 16x^3 + 9x^3 = ? (the exponent stays the same)

OpenStudy (anonymous):

@Vocaloid 25x^6

Vocaloid (vocaloid):

nope, remember what I said, the exponent stays the same

OpenStudy (anonymous):

Oh okay. Could you help me with a few more?

Vocaloid (vocaloid):

sure

OpenStudy (anonymous):

4. Find the difference. (9a^4 + 2a^2 + 5) - (16a^4 - 7a^2 + 1) A) 25a^4 - 5a^2 - 4 B) -7a^4 - 5a^2 + 6 C) 25a^8 + 9a^4 + 6 D) -7a^4 + 9a^2 + 4

Vocaloid (vocaloid):

start by combining like terms 9a^4-16a^4 = ?

OpenStudy (anonymous):

9a^6 ?

Vocaloid (vocaloid):

nope

Vocaloid (vocaloid):

9-16 = ?

OpenStudy (anonymous):

7

Vocaloid (vocaloid):

nope

OpenStudy (anonymous):

16-9 =7

Vocaloid (vocaloid):

right, but I asked 9-16, not 16-9

OpenStudy (anonymous):

9-16= -7

Vocaloid (vocaloid):

good! so, any ideas on what 9a^4 - 16a^4? (remember, when we add/subtract terms with exponents, the exponents don't change)

Vocaloid (vocaloid):

still there?

OpenStudy (anonymous):

@Vocaloid Hey, sorry I'm back now

Vocaloid (vocaloid):

good, so 9a^4 - 16a^4 = ?

OpenStudy (anonymous):

@Vocaloid -7a^4

Vocaloid (vocaloid):

good, now work on the a^2 term 2a^2 + 7a^2 = ?

OpenStudy (anonymous):

-7a^-4 + 9a^2 + 4

Vocaloid (vocaloid):

good!!